# BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis

### v1

@@ 6 Conclusion
We introduce BnSentMix, a novel sentiment analysis dataset tailored for code-mixed Bengali-English. Our work opens several potential research avenues for code-mixed Bengali. Researchers can explore other tasks, such as hate speech, offensive language, and abusive content detection on code-mixed data. Furthermore, the role of large language models and model ensembling techniques in processing code-mixed data can be further investigated. Our work addresses a significant gap for low-resource languages and sets a new standard for sentiment analysis in code-mixed Bengali-English.
@@
The label distribution of BnSentMix dataset is slightly imbalanced with only 9.2% samples labeled as mixed sentiment which can affect the performance of the model in classifying mixed sentiments. We also acknowledge that the sentiment of the annotator can be a source of bias during data annotation, though each data sample has been annotated twice by two different annotators, and annotation conflicts have been resolved by a data validator. Emojis and emoticons can convey sentiment but have been removed from our dataset. Their impact on the code-mixed sentiment can be explored in future works. Finally, complex sentiments involving sarcasm and humor are usually classified as neutral but can be subjective depending on the context.
@@
The hired data annotators were given monetary compensation which was significantly higher than the minimum wage of the region. Each annotator was given only $630$ data samples to annotate with no time restrictions. This ensured that the annotator did not overwork during data annotation. Annotator sentiment is subject to long working hours and can affect the sentiment labeling. To prevent this, we mandated five-minute breaks after every twenty-minute interval and provided refreshments upon request.

### v3

@@ 6 Conclusion
We introduce BnSentMix, a novel sentiment analysis dataset tailored for code-mixed Bengali-English. Our work opens several potential research avenues for code-mixed Bengali. Researchers can explore other tasks, such as hate speech, offensive language, and abusive content detection on code-mixed data. Our work addresses a significant gap for low-resource languages and sets a new standard for sentiment analysis in code-mixed Bengali-English.
@@
The label distribution of BnSentMix dataset is slightly imbalanced with only 9.2% samples labeled as mixed sentiment which can affect the performance of the model in classifying mixed sentiments. We also acknowledge that the sentiment of the annotator can be a source of bias during data annotation, though each data sample has been annotated twice by two different annotators, and annotation conflicts have been resolved by a third annotator.
@@
The hired data annotators were compensated significantly higher than the region’s minimum wage. Each annotator was only given around $630$ data samples with no time restrictions. This ensured that the annotator did not overwork during data annotation. Annotator sentiment is subject to long working hours and can affect sentiment labeling. To prevent this, we mandated five-minute breaks after every twenty-minute interval and provided refreshments upon request.
Acknowledgements
Our work is supported by the Islamic University of Technology Research Seed Grants (IUT RSG) (Ref: REASP/IUT-RSG/2022/OL/07/012). We sincerely appreciate Mohammed Saidul Islam and Md Mezbaur Rahman for guidance and Nejd Khadija for proofreading our work.